The sum of the measures of the interior angles of a convex polygon is 720degrees. The no. of sides of this polygon is?

thanks

Answer by Mathtut(3670)

(n-2)*180=720(sum of interiror angles) where n is the number of sides
:
n-2=4
:
n=6...so its a hexagon